# Document Type Classification using File Names

### v1

@@ Abstract
Rapid document classification is critical in several time-sensitive applications like digital forensics and large-scale media classification. Traditional approaches that rely on heavy-duty deep learning models fall short due to high inference times over vast input datasets and computational resources associated with analyzing whole documents. In this paper, we present a method using lightweight supervised learning models, combined with a TF-IDF feature extraction-based tokenization method, to accurately and efficiently classify documents based solely on file names that substantially reduces inference time. This approach can distinguish ambiguous file names from the indicative file names through confidence scores and through using a negative class representing ambiguous file names. Our results indicate that file name classifiers can process more than 80% of the in-scope data with 96.7% accuracy when tested on a dataset with a large portion of out-of-scope data with respect to the training dataset while being 442.43x faster than more complex models such as DiT. Our method offers a crucial solution for efficiently processing vast datasets in critical scenarios, enabling fast, more reliable document classification.
Document Type Classification using File Names
@@ 1 Introduction
Classifying massive amounts of documents at scale is often an important yet challenging task, especially at industrial scale, where applications of document classification include sensitive data identification, document review for legal cases, spam and malware flagging, etc. For example, in sensitive data identification, document sensitivity risk scores can be assigned according to type, e.g., academic_transcript $\rightarrow$ sensitive, medical_intake_form $\rightarrow$ sensitive, press_release $\rightarrow$ low-risk, etc.
Typical approaches to document type classification use text-, image-, layout-, and multimodal classifiers (e.g., Li et al. (2022); Huang et al. (2022); Pham et al. (2022)), but these models can be prohibitively slow at the scale required in many industry settings, as these models must execute sophisticated operations on the contents of a very large quantity of documents. Thus, there is a need for techniques that speed up the document classification pipeline to enable scale.
In this paper, we introduce and investigate the task of classifying documents by type (e.g., resume, press_release, etc.) based on their file names. Figure 1 displays several examples where documents can be categorized based on their file names. An added challenge is to correctly identify cases that are ambiguous. Since many file names are indicative of their contents (e.g., john_resume_2024.pdf $\rightarrow$ resume), we hypothesize that much simpler classifiers can be trained and used to categorize these documents.
Figure 1: Overview of the file name classification task. An ideal classifier can categorize indicative file names correctly, while also being able to defer predictions on ambiguous file names.
Our experimental result shows that the Random Forest file name classifier with a Trie tokenizer can achieve 99.6% accuracy on the indicative file names and 96.7% accuracy on a dataset distinct from the training dataset with a large portion of out-of-scope documents while deferring less than 80% of the in-scope documents. This Random Forest file name classifier requires only $1.23\times 10^{-4}$ seconds per prediction which is 442.43x faster than DiT Li et al. (2022), an exemplar image-based deep learning model that requires access to each document.
File Name Indicative or Ambiguous Type
@@
Intermolecular-USCAN.pdf ambiguous press_release
MO_Farmington.pdf ambiguous restaurant_menu
@@
Table 1: Examples of indicative file names and ambiguous file names.
@@ 2 Related Work
In this section we discuss prior work on classifying files using file names, and document type classification.
@@ 2.1 Classification with File Names
Document type classification via file names is an understudied topic. While prior research on categorizing content via file names has shown promise, it has largely focused on binary classification. For instance, Al-Nabki et al. (2023) proposed an approach to classify Child Sexual Exploitation Material (CSEM) through analyzing the file names and file paths of files when law enforcement officers seize a computer from a CSEM suspect due to the time-consuming nature of traditional and heavy weight models. Similarly, Jáñez Martino et al. (2021) accurately detected spam emails through analyzing email names using a Naive Bayes classifier, and Nguyen et al. (2019) demonstrated that a character-level CNN can accurately classify file names that contain enough information indicative of malware. Additionally, Sánchez-Paniagua et al. (2022) accurately detected phishing URLs using a logistic regression model combined with Token Frequency-Inverse Document Frequency (TF-IDF) feature extraction.
These studies demonstrate that binary classification through file names alone can be effective without deeper analysis of the contents of the files themselves. However, multi-class file name classification tasks, such as document type classification, is still under-researched. Moreover, the issue of handling file names that fall outside the the scope of the classifiers or those that lack clear indicators for accurate classification has not yet been adequately addressed in the existing literature.
2.2 Document Type Classification
Prior work on document type classification includes the development of large image-, text-, and page layout-based models for categorizing documents, mostly from datasets derived from the IIT-CDIP tobacco corpus (e.g., RVL-CDIP Harley et al. (2015) and Tobacco3482 Kumar and Doermann (2013)). These large models include the LayoutLM family (LayoutLM Xu et al. (2020), LayoutLMv2 Xu et al. (2021), LayoutLMv3 Huang et al. (2022)), DiT Li et al. (2022), DocFormer Appalaraju et al. (2021), Donut Kim et al. (2022), and LayoutLLM Fujitake (2024), etc. However, these models are costly to train and costly to run. Image-based models like CNNs and Transformers must operate on images of document pages, and text- and layout-based models must run on variants of the document from which text has been extracted via Optical Character Recognition (OCR), which adds a layer of complexity for each document that must be classified.
@@ 3 Problem Overview
In this study, we examine whether we can efficiently and accurately classify documents only based on their file names. Table 1 shows examples of file names and their categories. To enhance the clarity of our analysis, we distinguish the file names into the two following broad categories:
@@
  • Indicative: File names that contains the category name, a synonym of that category, or an abbreviation of the category name. Examples of indicative file names include john_smith_resume_2024.pdf (a resume document), john_smith_cv_2024.pdf (a resume document, where “cv” is used as a synonym for “resume”), and RTKRequestForm.pdf (a form document).
@@
  • Ambiguous: File names that does not contain the category name itself or a synonym of the category name. Examples of ambiguous file names include 20-005-10.pdf and document.pdf. We note that ambiguous file names are—to borrow terminology from Larson et al. (2019)—“out-of-scope” with respect to the in-scope or in-domain file type categories defined in the dataset.
@@
A key aspect of our approach is to determine whether the file name is indicative based on the confidence score associated with the predicted class for that file name. In cases where the confidence is too low, we defer the file name. In practice, we would instead use a more complex model at increased computational cost if the file is deferred by the filename classifier. In cases where the confidence score is high enough, we desire a highly accurate class prediction to forego the additional expense of running a deeper, more complex model involving the file’s contents. With a higher confidence threshold, the file name classifiers are more accurate with these predictions, but predicts a smaller portion of all the file names, ultimately adding computational burden to the heavier-duty model in the next stage. We discuss the trade-off and optimal confidence threshold for each file name classifier in Section 6. We also measure the AUROC of the confidence scores between indicative and ambiguous file names to analyze if the file name classifiers can produce confidence scores that differentiate between indicative and ambiguous file names.
@@
In this study, we only evaluate the inference speed of the heavier-duty models as we are interested how much faster the file name classifiers are compared to models such as DiT, rather than trying to outperform the more complex models in accuracy. This study aims to address whether file name classifiers can accurately predict indicative file names. Additionally, we evaluate if file name classifiers predict a large portion of the in-scope data with high accuracy when tested on a dataset containing indicative, ambiguous, and out-of-scope file names.
As shown in Li et al. (2024), adding a class of out-of-scope data in training improves classifier robustness when identifying out-of-scope data from the in-scope data. We hypothesize that by populating a new category with the ambiguous file names, coupled with marking predictions for supported categories with low confidence scores as ambiguous file names, the file name classifiers’ prediction accuracy can be improved.
@@ 4 Method
In this section, we discuss our approach to leveraging indicative and ambiguous file names to develop a fast, accurate document classification scheme that reduces the overall computation resources required to classify large volumes of documents. Our approach consists of a tokenization scheme for splitting input file names as well as training and evaluating light-weight machine learning pipeline.
@@ 4.1 Tokenization
We apply a Universal tokenizer to clean all input file names. Because many file names are gathered from URLs, we first each file name by replacing URL escape sequences (e.g., “%xx”) with the relevant single-character equivalent. We then split the file names by all punctuation characters, white space, and camel case boundaries (i.e., when a lower-case character is immediately followed by an upper-case character).
The Universal Tokenizer is not equipped to handle cases where file names contain words joined together without any punctuation or camel case such as "coursecatalog.pdf".
From randomly sampling 1,000 file names and annotating the Universal tokenizer’s performance, we discover that 3.4% file names needs further tokenization.
Figure 2 shows an example where the Universal tokenizer fails to separate “Coursesyllabus” into “Course” and “syllabus,” and additional processing is needed to correctly tokenize this file name.
To improve the tokenization on those file names, we first calculate the Token Frequency-Inverse Document Frequency of every token in relation to the category. After lemmatizing the output tokens with NLTK’s WordNetLemmatizer Bird and Loper (2004), tokens with TF-IDF higher than a threshold $k$ are identified as “keywords.” We choose $k$ to be 0.1 in this study. The trade-offs of selecting different value of $k$ is discussed in Section 6.5. The keywords are then added to a trie data structure to create a Trie tokenizer that can effectively locate and split important words that are joined together if the words in the file names are not separated with special characters, whitespace, or camel case. Another approach involves adding the "keywords" to a BERT Autotokenizer Wolf et al. (2020) using the “add_tokens” function in the Hugging Face library. For example, the Trie or BERT tokenizer will split file names such as “coursecatalog” in “course” and “catalog” to enable effective bag-of-words encoding.
@@ 4.2 Classification
Support Vector Machine (SVM), a supervised learning model that classifies data by finding a separating hyperplane that maximizes the margin. We use the linear kernel and the regularization parameter c = 0.1 as this combination yields the highest accuracy during our hyperparameter search where we compare the accuracy on the SVMs trained and tested on indicative file names with different kernels and value of c. The SVM model is implemented with Scikit-Learn which uses a one-vs-all approach.
Naive Bayes, a supervised learning model that learns the distribution of inputs by assuming that all the features are independent of each other. We implement this model with Scikit-Learn.
@@ 4.3 Metrics
For a selected confidence threshold $t$, we analyze the prediction accuracy by calculating #$\frac{\#\mathrm{correct}}{\#\mathrm{correct}+\#\mathrm{incorrect}}$, the number of correct predictions divided by the number of file names whose prediction confidence exceeds $t$. We examine the prediction rate by calculating #$\frac{\#\mathrm{correct}+\#\mathrm{incorrect}}{\mathrm{total}\#\mathrm{documents}}$, the sum of the Correct and Incorrect predictions over the total number of file names.
To analyze the models’ ability to distinguish ambiguous file names from the indicative file names, we calculate the confidence score for each prediction. For BERT, we use the highest softmax confidence score for each prediction Hendrycks and Gimpel (2016). For SVM, Naive Bayes, and Random Forest, we use the highest value from the “predict_proba” function through Scikit-Learn. Through calculating the Area Under the Receiver Operating Characteristic curve (AUROC) between the indicative file names and the ambiguous file names, we can measure how well the models can distinguish the ambiguous file names through confidence scores and avoid mispredicting file names that do not contain enough information. We note that AUROC has been used in prior work in the document classification space (e.g., Larson et al. (2022)).
@@ 5 Datasets
In this section we introduce two datasets used as part of our evaluation.
@@ 5.1 Web Search Dataset
We collected a dataset with 1,235 PDF files from 15 categories through web searches. A list of these 15 categories is shown in Table 6 in the Appendix.
We considered search prompts of the format: <category_name> filetype:pdf (where filetype:pdf specifies that we desire PDF documents from the web search engine). For example, we gather the PDF files for restaurant menus by searching restaurant menu filetype:pdf. Then, we manually open the files to verify that they belong to the desired category, and then record the verified document’s file name. We also note that synonyms or related words and phrases are also used when searching for category names. For instance, when searching for restaurant menus, we also searched for “diner menu,” “bar menu,” etc.
The Web Search dataset contains 15 categories, which is similar in number to the RVL-CDIP dataset Harley et al. (2015) which contains 16 categories. Table 1 displays more examples of indicative and ambiguous file names from this dataset. After manual annotation, we found that our Web Search dataset contains 978 indicative file names and 257 ambiguous file names. This breakdown of file names is used in our methodology to establish a “best-case” performance — for indicative file names, we desire a shallow, low-cost model that accurately classifies the given document into the best predicted category. The ambiguous file names that remain do not contain enough information to successfully classify, thus in practice we want the file name classifiers to defer to the heavier-cost model to analyze the content of those documents.
@@ 5.2 Common Crawl Dataset
In our evaluation, we also consider another open source dataset to demonstrate generalizability of our approach. We use a subset of the CC-MAIN-2021-31 dataset, which consists of PDFs, URLs, and metadata of eight million documents collected by Common Crawl from the web in July and August of 2021.222https://digitalcorpora.org/corpora/file-corpora/cc-main-2021-31-pdf-untruncated/ We randomly selected 1,198 documents that contain valid URLs and are in English. Of these 1,198 file names, there are 266 file names belonging to one of the 15 document type categories defined in the Web Search dataset from Section 5.1, and the remaining 932 file names belong to out-of-scope or categories unknown to the classifier. This distribution of categories for the file names differs from that of the Web Search dataset, and can simulate the robustness of file name classifiers trained on datasets that contain fewer categories and are smaller than the datasets they are tasked to predict when not every category is of interest to be classified.
The correct category of the Common Crawl dataset are obtained by manual annotation by two annotators without considering whether the file name is indicative or ambiguous. When the annotators cannot reach an agreement on the appropriate type, the document is discarded.
Models Without Negative Class With Negative Class
Tokenizer Threshold Accuracy Rate Tokenizer Threshold Accuracy Rate
SVM Trie 0.18 0.933 0.837 BERT 0.50 0.994 0.807
Naive Bayes BERT 0.15 0.893 0.877 BERT 0.22 0.979 0.803
Random Forest Trie 0.24 0.941 0.829 Trie 0.62 0.996 0.805
XGBoost Trie 0.13 0.898 0.895 Trie 0.60 0.993 0.809
BERT Trie 0.32 0.948 0.803 Trie 0.82 0.992 0.835
Table 2: The confidence threshold value and tokenizer combination that maximizes the prediction accuracy for every model trained with or without a negative class while the prediction rate is above 0.8 for all confidence threshold between 0 and 1 incremented by 0.01, trained and tested on the Web Search dataset.
6 Experiments

### v2

# Document Classification using File Names

@@ Abstract
Rapid document classification is critical in several time-sensitive applications like digital forensics and large-scale media classification. Traditional approaches that rely on heavy-duty deep learning models fall short due to high inference times over vast input datasets and computational resources associated with analyzing whole documents. In this paper, we present a method using lightweight supervised learning models, combined with a TF-IDF feature extraction-based tokenization method, to accurately and efficiently classify documents based solely on file names, that substantially reduces inference time. Our results indicate that file name classifiers can process more than 90% of in-scope documents with 99.63% and 96.57% accuracy when tested on two datasets, while being 442x faster than more complex models such as DiT. Our method offers a crucial solution to efficiently process vast document datasets in critical scenarios, enabling fast and more reliable document classification.
Document Classification using File Names
@@ 1 Introduction
Classifying massive amounts of documents at scale is often an important yet challenging task, especially at industrial scale, where applications of document classification include sensitive data identification, legal document review, spam and malware flagging, etc. Large-scale classification systems often process millions of documents, requiring robust models to handle diverse taxonomies efficiently Kandimalla et al. (2021). In sensitive data identification, document sensitivity risk scores can be assigned according to its category — for instance, academic transcripts and medical intake forms may be marked sensitive, while press releases may be marked low-risk.
Typical approaches to document classification use text-, image-, layout-, and multimodal classifiers (e.g., Li et al. (2022); Huang et al. (2022); Pham et al. (2022)), but these models can be prohibitively slow at the scale required in many industry settings Park et al. (2022), as these models must execute sophisticated operations on the contents of a very large quantity of documents. Thus, there is a need for techniques that speed up the document classification pipeline to enable scale.
In this paper, we introduce and investigate the task of classifying documents by categories (e.g., resume, press_release, etc.) using only their file names as input. Figure 1 displays several examples where documents can be categorized based on their file names. An added challenge is to correctly identify cases that are ambiguous, which are documents whose file names do not provide a clear indication of their content. However, since many file names are indicative of their contents (e.g., john_resume_2024.pdf $\rightarrow$ resume), we hypothesize that much simpler classifiers can be trained and used to categorize these documents. Furthermore, we hypothesize that a sufficiently large portion of documents can be classified using file name alone to yield a substantial savings in compute resources and time.
Figure 1: Overview of the file name classification task. An ideal classifier can categorize indicative file names correctly, while also being able to defer predictions on file names that lack enough category information.
Our experimental results show that the Random Forest file name classifier with a trie tokenizer achieves 99.63% and 96.57% accuracy on two respective datasets while deferring less than 10% of the in-scope documents. This Random Forest file name classifier requires only $1.23\times 10^{-4}$ seconds per prediction, which is 442x faster than DiT Li et al. (2022), an exemplar image-based deep learning model that requires access to each document’s content which contains both images and extracted text.
File Name Ambiguity Type Category
@@
privacy-notice.pdf out-of-scope privacy_notice
PremilinaryBalance_2011.pdf out-of-scope financial_sheet
Table 1: Examples of indicative file names, ambiguous, and out-of-scope file names.
@@ 2 Related Work
In this section, we discuss prior work on classifying files using file names, and document classification.
@@ 2.1 Classification with File Names
Document classification via file names is an understudied topic. While prior research on categorizing content via file names has shown promise, it has largely focused on binary classification. For instance, Al-Nabki et al. (2023) proposed an approach to classify Child Sexual Exploitation Material (CSEM) through analyzing the file names and file paths of files when law enforcement officers seize a computer from a CSEM suspect due to the time-consuming nature of heavy weight models. Similarly, Jáñez Martino et al. (2021) accurately detected spam emails through analyzing email names using a Naive Bayes classifier, and Nguyen et al. (2019) demonstrated that a character-level CNN can accurately classify file names that contain enough information indicative of malware. Additionally, Sánchez-Paniagua et al. (2022) accurately detected phishing URLs using a logistic regression model combined with Token Frequency-Inverse Document Frequency (TF-IDF) feature extraction.
These studies demonstrate that binary classification through file names or document titles alone can be effective without deeper analysis of the contents of the files themselves. However, multi-class file name classification tasks, such as document classification by categories, remains under-studied. Given the demand for efficient large-scale document processing in industry settings Eken et al. (2019), further research in this area is needed. Moreover, the issue of handling data that fall outside the scope of the classifiers or those that lack clear indicators for accurate classification illustrated in Larson et al. (2019) has not yet been adequately addressed in the existing literature for document classification.
2.2 Document Classification
Prior work on document classification includes the development of large image-, text-, and page layout-based models for categorizing documents, mostly from datasets derived from the IIT-CDIP tobacco corpus (e.g., RVL-CDIP Harley et al. (2015) and Tobacco3482 Kumar and Doermann (2013)). These large models include the LayoutLM family (LayoutLM Xu et al. (2020), LayoutLMv2 Xu et al. (2021), LayoutLMv3 Huang et al. (2022)), DiT Li et al. (2022), DocFormer Appalaraju et al. (2021), Donut Kim et al. (2022), and LayoutLLM Fujitake (2024), etc. However, these models are costly to train and costly to run. Image-based models like CNNs and Transformers must operate on images of document pages, and text- and layout-based models must run on variants of the document from which text has been extracted via Optical Character Recognition (OCR), which adds a layer of complexity for each document that must be classified.
@@ 3 Problem Overview
In this study, we examine whether we can efficiently and accurately classify documents only based on their file names. Table 1 shows examples of file names, their ambiguity types, and their categories. To enhance the clarity of our analysis, we distinguish the file names into the three following ambiguity types:
@@
  • Indicative: File names that contain the category name, a synonym of that category, or an abbreviation of the category name. Examples of indicative file names include john_smith_resume_2024.pdf (a resume document), john_smith_cv_2024.pdf (a resume document, where “cv” is used as a synonym for “resume”), and RTKRequestForm.pdf (a form document).
  • Ambiguous: File names that do not contain the category name itself or a synonym of the category name, but belongs to a category in the dataset. Examples of ambiguous file names include 20-005-10.pdf and document.pdf, which have names that are not indicative, but the content may happen to indicate the document belongs to a particular category in the dataset.
@@
  • Out-of-Scope: File names that do not belong to any category in the dataset. We note that out-of-scope file names are—to borrow terminology from Larson et al. (2019)—“out-of-scope” with respect to the in-scope or in-domain file categories defined in the dataset.
@@
A key aspect of our approach is to determine whether the file name is indicative based on the confidence score associated with the predicted class for that file name. In cases where the confidence is too low, we defer the file name. In practice, we would instead use a more complex model at increased computational cost if the file is deferred by the file name classifier. In cases where the confidence score is high enough, we desire a highly accurate class prediction to forego the additional expense of running a deeper, more complex model involving the file’s contents. With a higher confidence threshold, the file name classifiers are more accurate with these predictions, but predict a smaller portion of all the file names, ultimately adding computational burden to the heavier-duty model in the next stage. We discuss the trade-off and optimal confidence threshold for each file name classifier in Section 6. We also measure the AUROC of the confidence scores between indicative and ambiguous file names to analyze whether the file name classifiers produce confidence scores that differentiate between indicative and ambiguous or out-of-scope file names.
@@
In this study, we only compare the inference speed of the heavier-duty models as we are interested how much faster the file name classifiers are compared to models such as DiT, rather than trying to outperform the more complex models in accuracy. This study aims to address whether file name classifiers can accurately predict indicative file names, while deferring the ambiguous and out-of-scope file names. If a sufficiently large portion of documents contain indicative file names, then the savings of using a lightweight classifier before running a heavy duty model will substantially outweigh the cost of using a heavy duty model on the entire dataset.
@@ 4 Method
In this section, we discuss our approach to leveraging indicative and ambiguous file names to develop a fast, accurate document classification scheme that reduces the overall computational resources required to classify large volumes of documents. Our approach consists of a tokenization scheme for splitting input file names as well as training and evaluating a light-weight machine learning pipeline.
@@ 4.1 Tokenization
We apply a “universal tokenizer" to clean all input file names. The “universal tokenizer" first preprocess each file name by replacing escape sequences (e.g., “%xx”) with the relevant single-character equivalent. We then split the file names by all punctuation characters, white space, camel case boundaries (i.e., when a lower-case character is immediately followed by an upper-case character), and transitions between letters ([a-zA-Z]) and numbers ([0-9]).
The “universal tokenizer" is not equipped to handle cases where file names contain words joined together without any punctuation or camel case such as coursecatalog.pdf. From randomly sampling 1,000 file names and annotating the performance of the “universal tokenizer", we discover that 3.4% file names needs further tokenization. Figure 2 shows an example where the “universal tokenizer" fails to separate Coursesyllabus into “Course” and “syllabus,” instead requiring additional tokenization.
To improve the tokenization for such file names, we first calculate the Token Frequency-Inverse Document Frequency of every token in relation to the category. After lemmatizing the output tokens with NLTK version 3.6.7’s WordNetLemmatizer Bird and Loper (2004), tokens with TF-IDF higher than a threshold $k$ are identified as “keywords.” We choose $k$ to be 0.2 in this study. The trade-offs of selecting different value of $k$ is discussed in Section 6.6. The keywords are then added to a trie data structure to create a trie tokenizer that can effectively locate and split important words that are joined together if the words in the file names are not separated with special characters, whitespace, or camel case.
Another approach involves adding the “keywords” to a BERT tokenizer Wolf et al. (2020) using the “add_tokens” function in the Hugging Face library. For example, the trie or BERT tokenizer will split file names such as “coursecatalog” in “course” and “catalog” to enable effective bag-of-words encoding.
@@ 4.2 Classification
Support Vector Machine (SVM), a supervised learning model that classifies data by finding a separating hyperplane that maximizes the margin. We use the linear kernel and the regularization parameter $c=0.1$ as this combination yields the highest accuracy during our hyperparameter search where we compare the accuracy on the SVMs trained and tested on indicative file names with different kernels and value of $c$. The SVM model is implemented with Scikit-Learn version 1.4.2 Pedregosa et al. (2011) which uses a one-vs.-all approach.
Naive Bayes, a supervised learning model that learns the distribution of inputs by assuming that all the features are independent of each other. We implement this model with Scikit-Learn Pedregosa et al. (2011).
@@ 4.3 Metrics
For a selected confidence threshold $t$, we analyze the prediction accuracy by calculating #$\frac{\#\mathrm{correct}}{\#\mathrm{correct}+\#\mathrm{incorrect}}$, the number of correct predictions divided by the number of file names whose prediction confidence exceeds $t$. We examine the prediction rate by calculating #$\frac{\#\mathrm{correct}+\#\mathrm{incorrect}}{\mathrm{total}\#\mathrm{documents}}$, the sum of the correct and incorrect predictions over the total number of file names. Prediction Accuracy over 0.9 prediction rate is calculated to evaluate the models’ abilities to leverage confidence scores to avoid erroneous predictions while still processing the majority of the data.
Additionally, to analyze the models’ ability to distinguish ambiguous file names from the indicative file names, we calculate the confidence score for each prediction. For BERT, we use the highest softmax confidence score for each prediction Hendrycks and Gimpel (2016). For SVM, Naive Bayes, and Random Forest, we use the highest value from the predict_proba function through Scikit-Learn. Similar to Larson et al. (2022)’s work which used the Area Under Receiver Operating Characteristic curve (AUROC) to evaluate the models’ ability to distinguish out-of-scope samples from in-scope samples, we use AUROC to measure how well each model can distinguish ambiguous from indicative file names via confidence scores .
@@ 5 Datasets
We consider two datasets used as part of our evaluation.
@@ 5.1 Web Search Dataset
We collect a dataset with 1,500 PDF files with indicative file names from 15 categories through web searches, each category containing 100 files. A list of these 15 categories is shown in Table 9 in the Appendix.
We consider search prompts of the format: <category_name> filetype:pdf (where filetype:pdf specifies that we desire PDF documents from the web search engine). For example, we gather the PDF files for restaurant menus by searching restaurant menu filetype:pdf. To obtain file names, we select the substring after the last “/” from each document’s URL. Then, we manually open the files to verify that they belong to the correct category, and if their file names are indicative.
We also note that synonyms or related words and phrases are also used when searching for category names. For instance, when searching for restaurant menus, we also searched for “diner menu,” “bar menu,” etc.
The Web Search dataset contains 15 categories, which is similar in number to the RVL-CDIP dataset Harley et al. (2015) which contains 16 categories. Table 1 displays more examples of file names from this dataset.
We observe that numerous file names in the “resume” category contains people’s names. We replace the original last names by randomly generating last names using the Faker Python package version 36.1.1 joke2k (2016) for each of those file names.
@@ 5.2 Common Crawl Dataset
In our evaluation, we also consider another open source dataset to demonstrate generalizability of our approach. We use a subset of the CC-MAIN-2021-31 dataset, which consists of PDFs, URLs, and metadata of eight million unlabeled documents collected by Common Crawl from the web in July and August of 2021.222https://digitalcorpora.org/corpora/file-corpora/cc-main-2021-31-pdf-untruncated/ We randomly select 1,500 documents from the CC-MAIN-2021-31 dataset, and discard 302 documents that do not contain valid URLs, are not in English or results in label disagreements. Of these 1,198 file names, there are 284 indicative file names belonging to one of the 15 document categories defined in the Web Search dataset from Section 5.1, 264 ambiguous file names, and 650 out-of-scope file names.
The correct category labels and ambiguity types of the Common Crawl dataset are obtained by manual annotation by two annotators. When the annotators cannot reach an agreement on the appropriate category, the document is discarded.
6 Results